# Decentralized Parametric Stability Certificates for Grid-Forming Converter Control

### v5

@@ II. Loop Shifting
Given that $\mathcal{D}(s)$ and $\mathcal{N}(s)$ in Fig. 8(a) do not satisfy the passivity conditions in Theorem 1 (i.e., it can be shown that $\mathcal{N}(s)$ is not passive), we resort to dynamic loop-shifting techniques as presented in Section 2.3. More specifically, we consider the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ in Fig. 8(b) with the subsystems $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ given as
@@
We choose $\Gamma(s)=\mathrm{diag}(\Gamma_{1}(s),...,\Gamma_{n}(s))$ as a block-diagonal semi-stable transfer matrix with the $2\times 2$ matrix elements
@@
with the parameters
$\begin{split}\gamma_{1,i}^{\mathrm{p}}&=\gamma_{1,i}^{\mathrm{q}}=\tfrac{2|v|_{\mathrm{max}}^{2}}{\omega_{0}^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{2,i}^{\mathrm{p}}&=\gamma_{2,i}^{\mathrm{q}}=-3|v|_{\mathrm{max}}^{2}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{3,i}^{\mathrm{p}}&=3\tfrac{1}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}|v|_{\mathrm{max}}^{2}\\ \gamma_{3,i}^{\mathrm{q}}&=\tfrac{1}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\left(3|v|_{\mathrm{max}}^{2}+0.8\right),\end{split}$ (31)
whose structure loosely resembles the network model in 17.
@@ III. Passivity Checks & Theorem 1
For $\Gamma(s)$ selected as in 29, 30 and 31, we can show that $\mathcal{N}^{\prime}(s)$ is passive [^2] Notice that we can also show passivity of $\mathcal{N}^{\prime}(s)$ for the lossless case $\rho=0$ by evaluating 29, 30 and 31 at $\rho=0$ accordingly., i.e., it satisfies the conditions (i) to (iii) in Definition 2:
@@
$\begin{split}\mathcal{S}_{\mathcal{N}^{\prime}}(\mathrm{j}\omega)=\begin{bmatrix}\mathcal{S}_{\mathcal{N}^{\prime},11}(\mathrm{j}\omega)&\dots&\mathcal{S}_{\mathcal{N}^{\prime},1n}(\mathrm{j}\omega)\\ \vdots&\ddots&\vdots\\ \mathcal{S}_{\mathcal{N}^{\prime},n1}(\mathrm{j}\omega)&\dots&\mathcal{S}_{\mathcal{N}^{\prime},nn}(\mathrm{j}\omega)\end{bmatrix},\end{split}$ (32)
@@
$\begin{split}\mathcal{S}_{\mathcal{N}^{\prime},ii}(\mathrm{j}\omega)&=h_{\rho}(\omega)\left(\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\begin{bmatrix}-1&-\mathrm{j}\omega/\omega_{0}\\ \mathrm{j}\omega/\omega_{0}&-1\end{bmatrix}+\right.\\ &\quad\quad\quad\quad\quad\,\,\,\left.\begin{bmatrix}\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}&0\\ 0&\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}\end{bmatrix}\right)\\ \mathcal{S}_{\mathcal{N}^{\prime},ij}(\mathrm{j}\omega)&=h_{\rho}(\omega)|v|_{0,i}|v|_{0,j}b_{ij}\begin{bmatrix}1&\mathrm{j}\omega/\omega_{0}\\ -\mathrm{j}\omega/\omega_{0}&1\end{bmatrix}.\end{split}$ (33)
@@
$h_{\rho}(\omega)={4\rho\omega_{0}^{3}}/({(\omega_{0}^{2}+\rho^{2}\omega_{0}^{2}-\omega^{2})^{2}+4\rho^{2}\omega_{0}^{2}\omega^{2}})\geq 0.$ (34)
We can observe that 32 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries, i.e., the magnitude of the diagonal entry in a row is greater or equal to the sum of the magnitudes of the off-diagonal entries in that row. Namely, for the odd rows and $\forall\omega\geq 0$, we get
$\begin{split}|\textstyle\sum_{j\neq i}^{n}-|v|_{0,i}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}|&\geq\\ |\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}&|v|_{0,j}b_{ij}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}|v|_{0,j}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}-|v|_{\mathrm{max}}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}|&\geq\\ 2\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}+\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}\hskip 42.67912pt\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\geq\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(1+\tfrac{2\omega}{\omega_{0}}),\end{split}$
@@
$\begin{split}|\textstyle\sum_{j\neq i}^{n}-|v|_{0,i}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}|&\geq\\ |\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}&|v|_{0,j}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}|v|_{0,j}b_{ij}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}-|v|_{\mathrm{max}}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}|&\geq\\ 2\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}+\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\geq\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(1+\tfrac{2\omega}{\omega_{0}}).\end{split}$
@@
$\mathcal{D}(s)$$\mathcal{N}(s)$--$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$device dynamicsnetwork dynamics--$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
@@
$\mathcal{D}(s)$$\mathcal{D}^{\prime}$$\mathcal{N}(s)$$\mathcal{N}^{\prime}$--$\Gamma(s)$$\Gamma(s)$$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$device dynamicsnetwork dynamics$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$
@@
$\begin{split}\mathcal{R}_{\mathrm{j}0,ii}^{\mathcal{N}^{\prime}}&=\textstyle\sum_{j\neq i}^{n}b_{ij}\begin{bmatrix}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&0\\ 0&\tfrac{2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\end{bmatrix}+\\ &\quad\quad\quad\quad\quad\quad\quad\quad\,\,\,\begin{bmatrix}\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{p}}&0\\ 0&\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{q}}\end{bmatrix}\\ \mathcal{R}_{\mathrm{j}0,ij}^{\mathcal{N}^{\prime}}&=-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}.\end{split}$ (35)
Again, 35 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries. For the odd rows, we get
$\begin{split}|\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{p}}|&\geq\textstyle\sum_{j\neq i}^{n}|-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&\geq\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}.\end{split}$
For the even rows we get (with $|v|_{\mathrm{max}}=1.1$ and $|v|_{\mathrm{min}}=0.9$):
$\begin{split}|\textstyle\sum_{j\neq i}^{n}\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j})b_{ij}}{1+\rho^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{q}}|&\geq\textstyle\sum_{j\neq i}^{n}|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}+0.8)b_{ij}}{1+\rho^{2}}|&\geq\textstyle\sum_{j\neq i}^{n}|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{2|v|_{\mathrm{min}}^{2}-|v|_{\mathrm{max}}^{2}+0.8}{1+\rho^{2}}&\geq\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{\mathrm{max}^{2}}}{1+\rho^{2}}\end{split}$
Next, we derive the decentralized stability conditions in 24 and 25 under which $\mathcal{D}^{\prime}(s)$ is strictly passive. We consider $\mathcal{D}^{\prime}(s)=\text{diag}(\mathcal{D}_{1}^{\prime}(s),\dots,\mathcal{D}_{n}^{\prime}(s))$ with the matrix elements
$\begin{split}\mathcal{D}_{i}^{\prime}(s)=\mathcal{D}_{i}(s)(I-\Gamma_{i}(s)\mathcal{D}_{i}(s))^{-1}=\begin{bmatrix}\mathcal{D}_{\mathrm{p},i}^{\prime}(s)&0\\ 0&\mathcal{D}^{\prime}_{\mathrm{q},i}(s)\end{bmatrix},\end{split}$ (36)
@@
$\begin{split}\mathcal{D}_{\mathrm{q},i}^{\prime}(s)&=\tfrac{d_{\mathrm{q},i}s/|v|_{0,i}}{\tau_{\mathrm{q},i}s+1-\Gamma_{i}^{\mathrm{q}}(s)d_{\mathrm{q},i}s/|v|_{0,i}},\end{split}$ (37b)
for each of which we check the strict passivity conditions (i) and (ii) in Definition 3 independently. We start with $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$:
(i) Poles: To show that the poles of all elements of $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$ are in $\text{Re}(s)<0$, we start by inserting the expression for $\Gamma_{i}^{\mathrm{p}}(s)$ in 30 into 37a and obtain

### v6

@@ II. Loop Shifting
Given that $\mathcal{D}(s)$ and $\mathcal{N}(s)$ in Fig. 9(a) do not satisfy the passivity conditions in Theorem 1 (i.e., it can be shown that $\mathcal{N}(s)$ is not passive [24, 25]), we resort to dynamic loop-shifting techniques as presented in Section 2.3. More specifically, we consider the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ in Fig. 9(b) with the subsystems $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ given as
@@
Note again that loop-shifting with $\Gamma(s)$ is a mathematical tool for establishing passivity properties and does not involve any hardware or impedance modifications of the actual network or converter system. Having said that, we choose $\Gamma(s)=\mathrm{diag}(\Gamma_{1}(s),...,\Gamma_{n}(s))$ as a block-diagonal semi-stable transfer matrix with the $2\times 2$ matrix elements
@@
where the parameters are selected as
$\begin{split}\gamma_{1,i}^{\mathrm{p}}&=\gamma_{1,i}^{\mathrm{q}}=\tfrac{2|v|_{\mathrm{max}}^{2}}{\omega_{0}^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{2,i}^{\mathrm{p}}&=\gamma_{2,i}^{\mathrm{q}}=-3|v|_{\mathrm{max}}^{2}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{3,i}^{\mathrm{p}}&=-\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}\\ \gamma_{3,i}^{\mathrm{q}}&=-\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\tilde{\gamma}_{3,i}^{\mathrm{q}},\end{split}$ (32)
where $\tilde{\gamma}_{3,i}^{\mathrm{q}}\geq\tfrac{0.8}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}$. The choice of the transfer function $\Gamma(s)$ is guided by structural insights into the dynamic network model in 18, and algebraic considerations to ensure passivity of the loop-shifted system $\mathcal{N}^{\prime}(s)$, as detailed below.
[…]
Figure 8: Structural overview of the proof of Theorem 2.
@@ III. Passivity Checks & Theorem 1
For $\Gamma(s)$ selected as in 30, 31 and 32, we can show that $\mathcal{N}^{\prime}(s)$ is passive, i.e., it satisfies the conditions (i) to (iii) in Definition 2:
(i) Poles: The poles of all elements of $\mathcal{N}^{\prime}(s)$ are $p_{1}=\mathrm{j}0$ and $p_{2,3}=-\rho\omega_{0}\pm\mathrm{j}\omega_{0}$, i.e., $\text{Re}(p_{k})\leq 0$ for $k\in\{1,2,3\}$.
@@
[…]
@@
$\hskip-5.69054pth_{\rho}(\omega)={4\rho\omega_{0}^{3}}/({(\omega_{0}^{2}\hskip-1.42262pt+\hskip-1.42262pt\rho^{2}\omega_{0}^{2}\hskip-1.42262pt-\hskip-1.42262pt\omega^{2})^{2}\hskip-1.42262pt+\hskip-1.42262pt4\rho^{2}\omega_{0}^{2}\omega^{2}})\geq 0.$ (35)
We can observe that 33 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries, i.e., the magnitude of the diagonal entry in a row is greater or equal to the sum of the magnitudes of the off-diagonal entries in that row. Namely, for the odd rows and $\forall\omega\geq 0$, we get
$\begin{split}|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{0,i}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{p}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{p}}|\hskip-1.70717pt&\geq\\ |\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{0,i}^{2}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|\hskip-0.85358pt|v|_{0,i}&|v|_{0,j}b_{ij}|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|\hskip-0.85358pt|v|_{0,i}|v|_{0,j}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|\hskip-5.69054pt\\ \Leftrightarrow\hskip-1.42262pt|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{\mathrm{max}}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{p}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{p}}|\hskip-1.70717pt&\geq\hskip-1.70717pt\\ 2\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\hskip-1.42262pt|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\hskip-1.42262pt|v|_{\mathrm{max}}^{2}b_{ij}\hskip 42.67912pt\\ \Leftrightarrow\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})\hskip-1.70717pt&\geq\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}\hskip-0.56905pt(1+\tfrac{2\omega}{\omega_{0}}),\end{split}$
@@
$\begin{split}\hskip-5.69054pt|\hskip-0.85358pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{0,i}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{q}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\\ |\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{0,i}^{2}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|\hskip-0.85358pt|v|_{0,i}&|v|_{0,j}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}\hskip-0.85358pt|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|\hskip-0.85358pt|v|_{0,i}|v|_{0,j}b_{ij}\hskip-0.85358pt|\hskip-1.99168pt\hskip-5.69054pt\\ \hskip-5.69054pt\Leftrightarrow|\hskip-0.85358pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{\mathrm{max}}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{q}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\\ 2\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}\\ \hskip-5.69054pt\Leftrightarrow\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}\hskip-0.56905pt(1+\tfrac{2\omega}{\omega_{0}}).\end{split}$
@@
(iii) Imaginary poles: For $\rho\neq 0$, $\mathcal{N}^{\prime}(s)$ has one imaginary pole, i.e., $p_{1}=\mathrm{j}0$, which is a simple pole. We therefore compute the limit $\mathcal{R}_{\mathrm{j}0}^{\mathcal{N}^{\prime}}\coloneqq\text{lim}_{s\rightarrow\mathrm{j}0}(s-\mathrm{j}0)\mathcal{N}^{\prime}(s)$, where each $\mathcal{R}_{\mathrm{j}0,ij}^{\mathcal{N}^{\prime}}$ represents a $2\times 2$ transfer matrix block. The diagonal and off-diagonal elements are given by
[…]
Again, 36 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries. For the odd rows, we get
$\begin{split}\hskip-5.69054pt|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\gamma_{3,i}^{\mathrm{p}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|\hskip-0.85358pt-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\hskip-1.99168pt\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}.\end{split}$
For the even rows we get (with $|v|_{\mathrm{max}}\hskip-1.42262pt=\hskip-1.42262pt1.1$ and $|v|_{\mathrm{min}}\hskip-1.42262pt=\hskip-1.42262pt0.9$):
$\begin{split}\hskip-5.69054pt|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j})b_{ij}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\gamma_{3,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\hskip-1.99168pt\\ \Leftrightarrow|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j})b_{ij}}{1+\rho^{2}}+\tilde{\gamma}_{3,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\hskip-1.99168pt\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{2|v|_{\mathrm{min}}^{2}-|v|_{\mathrm{max}}^{2}}{1+\rho^{2}}+\tilde{\gamma}_{3,i}^{\mathrm{q}}&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt\hskip-0.85358ptb_{ij}\tfrac{|v|_{\mathrm{max}^{2}}}{1+\rho^{2}}\hskip-1.99168pt\end{split}$
$\mathcal{D}(s)$$\mathcal{N}(s)$-$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$device dynamicsnetwork dynamics-$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
@@
$\mathcal{D}(s)$$\mathcal{D}^{\prime}$$\mathcal{N}(s)$$\mathcal{N}^{\prime}$-$\Gamma(s)$$\Gamma(s)$$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$device dynamicsnetwork dynamics[$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$
@@
Next, we derive the decentralized stability conditions in 25 and 26 under which $\mathcal{D}^{\prime}(s)$ is strictly passive. We consider $\mathcal{D}^{\prime}(s)=\text{diag}(\mathcal{D}_{1}^{\prime}(s),\dots,\mathcal{D}_{n}^{\prime}(s))$ with the matrix elements
@@
$\begin{split}\mathcal{D}_{\mathrm{q},i}^{\prime}(s)&=\tfrac{\tfrac{d_{\mathrm{q},i}s}{|v|_{0,i}}}{\tau_{\mathrm{q},i}s+1-\Gamma_{i}^{\mathrm{q}}(s)\tfrac{d_{\mathrm{q},i}s}{|v|_{0,i}}},\end{split}$ (38b)
for each of which we check the strict passivity conditions (i) and (ii) in Definition 3 independently. We start with $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$ and insert the expression for $\Gamma_{i}^{\mathrm{p}}(s)$ in 31 into 38a, i.e.,